# Effective Note-taking and its Impact on Learning Undergraduate Introductory Physics Courses

### v1

@@ Abstract
Taking notes during lectures is one of the required skills, among many others, that students need (i) to master the topic covered in the lecture, (ii) to actively engage in the learning process with minimal to no distractions, (iii) to retain learned knowledge and skills for a longer time, and (iv) in securing better grade. To learn the role of note-taking in learning undergraduate-level introductory algebra-based physics courses, we present a comparative study of students’ achievement in two minority-serving universities, namely, Fayetteville State University and The University of North Carolina at Pembroke (i) when students were taught effective note-taking strategies, motivated them to prepare notes and let them use their self-prepared notes in timed assignments and exams versus (ii) no note-taking scheme was implemented, keeping all other conditions the same. The no-note-taking scheme was used in the first introductory algebra-based physics course, while the note-taking scheme was used in the second introductory algebra-based physics course. Students’ scores in the final exam and aggregated grade accounting for all grading components are taken as measurement tools. Analysis shows that effective note-taking helps students secure better letter grades.
@@ I Introduction
An effective learning process requires good organization of learning materials and attentiveness while learning. Students cannot sustain long-term attention to learning materials if activities that allow them to actively engage and contribute are omitted. In a typical lecture with no activities, students’ attentiveness peaks ( 70%) at around 10 minutes and declines after that, reaching a minimum ( 20%) at around 20 minutes [3, 4]. To maintain students’ attentiveness at the maximum possible level throughout the lecture, an instructor can incorporate various activities, visuals, simulations, and other engaging elements at regular intervals. Active processing helps move information from working memory to long-term memory [5], and taking notes is one way to do this. In this study, we explore how learning can be improved by encouraging effective note-taking and adding activities during lectures.
@@
Effects of note-taking on learning a secondary language have been extensively studied, with the literature showing that it helps in learning the second language (see a recent review in Ref. [6]). In learning the second language vocabulary, Jin and Webb reported that vocabulary learning gains in the range of 12.5% to 23.0% [7]. Authors in Ref. [8] presented that note-taking has even better impacts on second language listening comprehension for high school students. Campana conducted a study  [9] on the effectiveness of using guided notes in high school physics by dividing the class into two groups. Therein, the author observed that the sections that were completed and used guided notes performed better academically.
@@
To the best of our knowledge, no study has examined the role of effective note-taking in learning algebra-based undergraduate physics courses, as we presented in this research work to compare our findings; however, some works in the related science, technology, engineering, and mathematics (STEM) fields are available. Anastasiou et al. studied the effectiveness of concept maps in science learning [10]. Therein the authors used data from 55 studies published between 1980 and 2020, including peer-reviewed journals and dissertations which covered various teaching settings, conditions, and methods and divulged that the concept maps were associated with increased science learning, but there was a significant heterogeneity in most subsets of data. In one of the studies made by Shi et al on 42 college students’ learning achievement and cognitive load in a 6-week lecture-based computer network course, the authors found that note taking helped in learning   [11]. In this regard, the authors demonstrated that among the three different note-taking styles observed, namely collaborative note-taking, laptop note-taking, and traditional longhand note-taking, collaborative note-taking was found to improve learning achievement. Recently, Fanguy studied the effects of collaboratively generated notes by students versus instructor-provided notes on recalling and academic writing performance and found that student generated collaborative notes by students were more effective inenhancing academic writing, as well as in applying and developing related skills [12].
@@
With the current technological advancement in the electronic industry, students are likely to suffer from many digital distractions in the learning environment, which interfere badly with learning [13]. In search of a better solution to minimize distraction in learning and increase attentiveness, we here present a case study on the role of effective note-taking in learning undergraduate physics courses in two minority-serving institutions Fayetteville State University (FSU) and The University of North Carolina at Pembroke (UNCP) located in the Sandhills region of North Carolina.
@@ II Effective Note-taking
Note-taking significantly enhances deeper understanding, leading to metacognition, as writing notes helps the brain actively process and encode information, link ideas, integrate them with new knowledge, strengthen memory pathways, provide opportunities for reviewing, self-testing, and encourage reflection [14]. Note-taking offers students an elaborative rehearsal, supports analytical and critical thinking, and facilitates repetition and paraphrasing of the information, thereby enhancing their long-term memory [15]. Training students in note-taking skills, encouraging them to explore further, and helping them become effective notetakers are crucial. Taking an example of an undergraduate algebra-based physics course, the survey was conducted based on the hypothesis that note-taking helps in the learning process.
@@
Note-taking is a crucial aspect of effective learning as it forces learners to engage with the learning material actively. It helps reinforce understanding and retention, encode information into long-term memory, provide opportunities for repetition and rehearsal, and organize information in a structured manner, making it easier to follow the logical flow of concepts. The preferred way of learning or approaching a task varies from person to person based on how the mind receives and processes information from its senses. Learners’ productivity, creativity, problem-solving skills, decision-making capability, quality of learning, and achievement can be increased and speed up by properly recognizing and understanding the learners’ preferred learning style [16]. Notes can be personalized to fit an individual’s learning styles, integrating mnemonics, incorporating annotations and highlights that give the full picture, making a good sense of the concepts to the note-taker [17]. Well-prepared notes provide opportunities for review, revision, and repetition, essential for memorizing formulas, principles, and problem-solving techniques, improving memory [18].
Numerous note-taking methods are in practice to make the learning process easy, effective, and time-saving. Unfortunately, there is no single method of note-taking that can serve the best for all. Different note-taking methods, their effectiveness, and taking the best notes have been the center of research for quite a long time. For example, see Refs. [31, 28, 25, 19, 32, 27, 20, 22, 21, 24, 23, 26, 29, 30].
A brief review of different note-taking methods is in order. Cornell note-taking is one of the most commonly used techniques of taking notes, in which each page is divided into four segments, namely, a line at the top for the topic, a few lines at the bottom for the summary, and two columns divided for the middle part of the paper with a narrow left for key points and wider right column for an extensive detail of the key corresponding points [33]. Significantly structured notes with main topic(s) followed by indented subtopic(s), sub-subtopic(s), and detailed elaborations underneath in a logical manner forming a skeleton of the textbook chapter, called outline method, serves as an excellent study guide when preparing for tests  [34]. One can also use the boxing method of note-taking, which involves drawing boxes around related pieces of information to visually organize and categorize thoughts, concepts, or ideas of the same category by visually separating different topics under different boxes and titles, making it easier to review and memorize the notes for longer [35]. The Zettelkasten method of note-taking is a personal knowledge management system, which serves well in creating an ever-growing web of knowledge over a certain topic, connecting fragmented note cards having one idea per note to a structured storing box, helping in the accumulation of knowledge on a specific topic over time [36]. In addition to being a memory retention tool, the Zettelkasten method fosters a deeper understanding of complex subjects and supports generating new ideas and insights, boosting creativity, productivity, and adaptability in organizing and processing information.
@@
Figure 1: The figure shows an illustration of the mind mapping method about ways how students can increase attentiveness in the classroom to get a better learning experience.
@@
One of the most sophisticated, dynamic, and visually stimulating note-taking methods, in which a learner writes the main idea in the center of the page and branches out with connecting lines to subtopics, supporting details, and related concepts, creating connections between them, is the mind mapping method [37]. Figure 1 illustrates a mind map of strategies that students can follow to increase attentiveness in the classroom. Some key strategies that students need to embrace to improve attentiveness are taking a digital detox, minimizing digital distractions, taking effective notes, actively participating in class, learning autonomy, properly managing tasks, and following a healthy lifestyle. The central concept of increasing attentiveness sits at the center and branches outwards to the six listed strategies, as shown in Fig. 1. The secondary, tertiary, and quaternary branches can be created, extending the key point to add sub-subtopics and details. Each strategy in Fig. 1 further branches outwards, connecting different approaches students may want to follow. For example, the note-taking strategy connects tactics of having handwritten notes instead of digital or online application written notes, adding visual concepts, using memory aids, preparing structurally consistent notes, and having mnemonics, symbols, and abbreviations instead of writing complete sentences or paragraphs.
@@
A digital detox helps reduce stress by alleviating cognitive overload, improve memory, enhance focus, and allow the brain to rest and process information more effectively, thereby improving mental clarity [38]. Moreover, the digital detox helps in students’ psychological development and improvement of students’ concentration [39]. Although the use of digital technology is unavoidable in the current digital transformation era, a brief discussion on students’ necessity of taking a digital detox during classes to increase attentiveness is in order. Unless the learning strategy, instructional approach, and subject and content-like contextual factors demand, a wise choice for the students to improve attentiveness in class is not to use network and internet-connected electronic devices such as cell phones, smartwatches, laptops, tablets, etc., during lectures, which not only distract students but also negatively impact in students’ participation in classroom activities. Students are strongly recommended to turn off notifications during class and follow daily communication management strategies such as setting disconnection time, no screen time, communication preferred time, time for social media interaction, and decreasing internet addiction, etc. Mind mapping diagrams are free-flowing and helpful for brainstorming, finding linkage between ideas/concepts/topics, and generating ideas, keeping the main topic at the center. One may choose to prepare or recommend students to create concept maps as they are more structured, focused on connecting different concepts, and follow a hierarchy. Taking data from 55 studies about concept mapping in science achievement published in peer-reviewed journals and dissertations between 1980 and 2020, Anastasiou et al analyzed the effectiveness of concept maps science learning across several instructional settings, conditions, and methodological features [10]. Therein, the authors found that the concept maps were associated with increased science learning but there was a significant heterogeneity in most subsets of data.
@@
Many other effective note-taking methods are also in practice such as the charting method, outline method, sentence method, flow method, boxing method, question-evidence-conclusion (QEC) method, split page method, writing on slides, bullet point method, list method, and rapid logging method [40, 41, 42, 43, 44, 45, 46]. The choice of note-taking method depends on many factors such as learning style, subject matter and topic, lecture structure and content delivery method, context, learner’s personal goal and preferences, available time to prepare notes, and need for mathematical, logical, and visual sophistication.
@@ III.1 Class and Lecture Design
The role of note-taking in learning introductory algebra-based undergraduate physics courses was studied in the minority-serving institutions, FSU and UNCP. The courses were designed based on the backward design of effective teaching, prioritizing the intended learning outcomes and course objectives over the chapters to be covered in the course, as well as the choice of instructional methods and assessment types.
@@
A brief discussion on the backward design of instructional method is in order. The backward design instructional method is based on aligning instruction, course content, pedagogy, and assessment with the learning objective, in contrast to the traditional teaching method of starting the learning process based on teacher-designed content, delivering lectures, assigning problems, and assessing students based on what was taught [47]. In the backward design method, one first determines the desired learning objectives and goals that students are expected to learn by the end of the day, chapter, and the course. Second, assessments are determined and planned to check whether students have learned what they were expected to learn. Third, learning activities and tasks are prepared that will be done in the classroom and beyond to prepare students for assessments. Algebra-based physics is primarily designed for students majoring in life sciences rather than students majoring in physical science. Students pursuing careers in life sciences are often less motivated and less prepared for physics courses, which require active, goal-oriented learning rather than passive, teacher-centered methods. To address this, we used a backward design approach in teaching these courses.
@@
The classes met twice a week for 75 minutes each day. Pedagogy includes lecturing, simulations, demonstrations, problem-solving, hands-on activities, think-pair-share, and class work as multiple choice questions. The classwork covered topics discussed at the beginning and end of the lecture, which helped ensure that students arrived on time and stayed for the whole class and boosted class attendance [13]. Students were strongly recommended to take a digital detox in the classroom, which minimizes digital distractions and helps minimize multi-tasking, which is considered one of the major learning pitfalls [48]. Students using electronic devices to take notes are prone to multitasking, but multitasking during learning impacts comprehension and knowledge retention as attention gets divided, focus decreases, the brain needs to pay the task-switching costs on refocusing, the brain receives cognitive load usually more than it can handle with no stress, and working memory is disrupted [49]. The assignments include online quizzes, mostly focused on concept development, and instructor-prepared homework assignments focused on problem-solving. The instructors used a liquid syllabus, which is a dynamic, web-based, interactive, and student-engaging syllabus that includes visual elements such as figures and videos, links to resources, and QR codes. The liquid syllabus has fixability in terms of dates, learning activities, and so on, to ensure the course objective is met by the end of the semester. The liquid syllabus, featuring more multimedia components, was adopted, inspired by the fact that students rarely read the syllabus once at the beginning of the semester and need to be made aware of the instructors’ expectations, requirements, and core objectives for the content and the course. Study shows liquid syllabus provides students with what they need to succeed in the course, increases equity, gives a sense of belonging and respect using student-friendly language and visuals, and provides students with a humanizing introduction to the course [50].
Students must have one prerequisite course to enroll in Physics-I at FSU and UNCP. At FSU, students must have a minimum passing grade in one math course out of 4 math courses college trigonometry, algebra and trigonometry, pre-calculus mathematics-I, or pre-calculus mathematics-II. At UNCP, students must have a minimum passing grade in one math course out of college algebra, pre-calculus, or calculus -I or II. Comparison of course contents and the academic standards show that the minimum criterion of the prerequisite for Physics-I is the same. In both institutions, successful completion of Physics-I is the prerequisite for Physics-II. To be noted, the general Physics course at FSU is equivalent to the college Physics course at UNCP.
III.2 Grading Policy
@@
The grading practices (followed by Tables 1 and 2) were communicated at the beginning of the semester and frequently during the semester, and grade points are made student-trackable to ensure students understand the grading practices, can determine their weaknesses and strengths, and are focused on their goals. Assignments were structured to (i) reduce the negative impact of late work, keeping students responsible for submitting assignments on time, (ii) give chances to implement feedback for improvement, and (iii) provide opportunities for students to reflect on and improve their performance, which motivated student in learning by doing and revising. We focused on reducing the basis of grading by concentrating on academic achievement and excellence by excluding non-academic components such as attendance, seminar participation, group discussion, extra work for points, etc. Online tests were provided with three possible attempts for each question, in which a link to the relevant section in the eTextbook and available resources was provided. A hint with insight into how to answer the questions was made visible after the first attempt, and the detailed solutions were provided a minute after the deadline. Numerical values were randomized for each student, and the order of answer choices was shuffled in each attempt. The desired number of significant figures was made visible for each question. Each student got feedback on each attempt. The best score out of the maximum three attempts was recorded.
One of the bonafide meanings of learning is storing the received knowledge, concepts, and gained skills in the long-term memory, retrieving them effortlessly as needed, and applying them in relevant contexts. Senses first perceive knowledge, information, course contents, or any other learning materials, pass to the sensory memory, go to the working memory, and finally, it either gets lost or gets stored in the long-term memory based on the learner’s attention, elaborated rehearsal, and repetition [3, 4]. Students are encouraged to increase attentiveness, as we presented in Fig 1, and be active physically and mentally in the learning environment rather than just being passive listeners. Recognizing that elaborated rehearsal, implementation of feedback, repetition of tasks, and gaining overall skills are crucial, we assigned a lower weight to problem-solving homework assignments and online quizzes than terminal exams and the final comprehensive exam. For the same reason, the final exam had a slightly larger weight than each terminal exam.

### v3

@@ Abstract
Taking notes during lectures is one of the required skills, among many others, that students need (i) to master the topic covered in the lecture, (ii) to actively engage in the learning process with minimal to no distractions, (iii) to retain learned knowledge and skills for a longer time, and (iv) in securing higher letter grades. To learn the role of notetaking in learning undergraduate-level introductory physics courses, we present a comparative study of students’ achievement in a mid-terminal exam at a historically black Fayetteville State University (FSU) (i) when students were taught effective notetaking strategies, motivated them to prepare notes and let them use their self-prepared notes in a terminal exam versus (ii) no notetaking scheme was implemented, keeping all other conditions the same. The no-notetaking scheme was used in 4 different sections over a few semesters, and the notetaking scheme was used at the same level of an introductory physics course in 3 different sections in other semesters. Students’ scores in one of the mid-term exams are taken as measurement tools. Grade analysis indicates that effective notetaking enhances students’ letter grades and lowers failure rates.
@@ I Introduction
One of the bonafide meanings of learning is storing the received knowledge, concepts, and gained skills in the long-term memory, retrieving them effortlessly as needed, and applying them in relevant contexts. Senses first perceive knowledge, information, course contents, or any other learning materials, pass to the sensory memory, go to the working memory, and finally, it either gets lost or gets stored in the long-term memory based on the learner’s attention, elaborated rehearsal, and repetition [3, 4]. An effective learning process requires good organization of learning materials and attentiveness while learning. Students cannot sustain long-term attention to learning materials if activities that allow them to actively engage and contribute are omitted. In a typical lecture with no activities, students’ attentiveness peaks (70%) at around 10 minutes and declines after that, reaching a minimum (20%) at around 20 minutes [3, 4]. To maintain students’ attentiveness at the maximum possible level throughout the lecture, an instructor can incorporate various activities, visuals, simulations, and other engaging elements at regular intervals. Active processing helps move information from working memory to long-term memory [5], and taking notes is one way to do this. In this study, we explore how learning can be improved by encouraging effective notetaking and adding activities during lectures.
@@
Effects of notetaking on learning a secondary language have been extensively studied, with the literature showing that it helps in learning the second language (see a recent review in Ref. [6]). In learning the second language vocabulary, Jin and Webb reported that vocabulary learning gains in the range of 12.5% to 23.0% [7]. Authors in Ref. [8] presented that notetaking has even better impacts on second language listening comprehension for high school students. Campana conducted a study  [9] on the effectiveness of using guided notes in high school physics by dividing the class into two groups. Therein, the author observed that the sections that were completed and used guided notes performed better academically.
@@
To the best of our knowledge, no study has examined the role of effective notetaking in learning algebra-based undergraduate physics courses, as we presented in this research work to compare our findings; however, some works in the related science, technology, engineering, and mathematics (STEM) fields are available. Anastasiou et al. studied the effectiveness of concept maps in science learning [10]. Therein the authors used data from 55 studies published between 1980 and 2020, including peer-reviewed journals and dissertations which covered various teaching settings, conditions, and methods and divulged that the concept maps were associated with increased science learning, but there was a significant heterogeneity in most subsets of data. In one of the studies made by Shi et al on 42 college students’ learning achievement and cognitive load in a 6-week lecture-based computer network course, the authors found that note taking helped in learning [11]. In this regard, the authors demonstrated that among the three different notetaking styles observed, namely collaborative notetaking, laptop notetaking, and traditional longhand notetaking, collaborative notetaking was found to improve learning achievement. Recently, Fanguy studied the effects of collaboratively generated notes by students versus instructor-provided notes on recalling and academic writing performance and found that student generated collaborative notes by students were more effective inenhancing academic writing, as well as in applying and developing related skills [12].
@@
With the current technological advancement in the electronic industry, students are likely to suffer from many digital distractions in the learning environment, which interfere badly with learning [13]. In search of a better solution to minimize distraction in learning and increase attentiveness, we here present a case study on the role of effective notetaking in learning introductory undergraduate physics courses.
@@ II Effective Note-taking
Notetaking significantly enhances deeper understanding, leading to metacognition, as writing notes helps the brain actively process and encode information, link ideas, integrate them with new knowledge, strengthen memory pathways, provide opportunities for reviewing, self-testing, and encourage reflection [14]. Notetaking offers students an elaborative rehearsal, supports analytical and critical thinking, and facilitates repetition and paraphrasing of the information, thereby enhancing their long-term memory [15]. Training students in notetaking skills, encouraging them to explore further, and helping them become effective notetakers are crucial. Taking an example of an undergraduate algebra-based physics course, the survey was conducted based on the hypothesis that notetaking helps in the learning process.
@@
Notetaking is a crucial aspect of effective learning as it forces learners to engage with the learning material actively. It helps reinforce understanding and retention, encode information into long-term memory, provide opportunities for repetition and rehearsal, and organize information in a structured manner, making it easier to follow the logical flow of concepts. The preferred way of learning or approaching a task varies from person to person based on how the mind receives and processes information from its senses. Learners’ productivity, creativity, problem-solving skills, decision-making capability, quality of learning, and achievement can be increased and speed up by properly recognizing and understanding the learners’ preferred learning style [16]. Notes can be personalized to fit an individual’s learning styles, integrating mnemonics, incorporating annotations and highlights that give the full picture, making a good sense of the concepts to the note-taker [17]. Well-prepared notes provide opportunities for review, revision, and repetition, essential for memorizing formulas, principles, and problem-solving techniques, improving memory [18].
@@
Figure 1: The figure is an illustration of the mind-mapping method, showing ways to increase attentiveness in the classroom to improve the learning experience for students.
Numerous notetaking methods are in practice to make the learning process easy, effective, and time-saving allowing students to choose the one that works best for them. Different notetaking methods, their effectiveness, and taking the best notes have been the center of research for quite a long time. For example, see Refs.[19-32] .
A brief review of different notetaking methods is in order. Cornell notetaking is one of the most commonly used techniques of taking notes, in which each page is divided into four segments, namely, a line at the top for the topic, a few lines at the bottom for the summary, and two columns divided for the middle part of the paper with a narrow left for key points and wider right column for an extensive detail of the key corresponding points [33]. Significantly structured notes with main topic(s) followed by indented subtopic(s), sub-subtopic(s), and detailed elaborations underneath in a logical manner forming a skeleton of the textbook chapter, called outline method, serves as an excellent study guide when preparing for tests  [34]. One can also use the boxing method of notetaking, which involves drawing boxes around related pieces of information to visually organize and categorize thoughts, concepts, or ideas of the same category by visually separating different topics under different boxes and titles, making it easier to review and memorize the notes for longer [35]. The Zettelkasten method of notetaking is a personal knowledge management system, which serves well in creating an ever-growing web of knowledge over a certain topic, connecting fragmented note cards having one idea per note to a structured storing box, helping in the accumulation of knowledge on a specific topic over time [36]. In addition to being a memory retention tool, the Zettelkasten method fosters a deeper understanding of complex subjects and supports generating new ideas and insights, boosting creativity, productivity, and adaptability in organizing and processing information.
@@
One of the most sophisticated, dynamic, and visually stimulating notetaking methods, in which a learner writes the main idea in the center of the page and branches out with connecting lines to subtopics, supporting details, and related concepts, creating connections between them, is the mind mapping method [37]. Figure 1 illustrates a mind map of strategies that students can follow to increase attentiveness in the classroom. Some key strategies that students need to embrace to improve attentiveness are taking a digital detox, minimizing digital distractions, taking effective notes, actively participating in class, learning autonomy, properly managing tasks, and following a healthy lifestyle. The central concept of increasing attentiveness sits at the center and branches outwards to the six listed strategies, as shown in Fig. 1. The secondary, tertiary, and quaternary branches can be created, extending the key point to add sub-subtopics and details. Each strategy in Fig. 1 further branches outwards, connecting different approaches students may want to follow. For example, the notetaking strategy connects tactics of having handwritten notes instead of digital or online application written notes, adding visual concepts, using memory aids, preparing structurally consistent notes, and having mnemonics, symbols, and abbreviations instead of writing complete sentences or paragraphs.
@@
A digital detox helps reduce stress by alleviating cognitive overload, improve memory, enhance focus, and allow the brain to rest and process information more effectively, thereby improving mental clarity [38]. Moreover, the digital detox helps in students’ psychological development and improvement of students’ concentration [39]. Although the use of digital technology is unavoidable in the current digital transformation era, a brief discussion on students’ necessity of taking a digital detox during classes to increase attentiveness is in order. Unless the learning strategy, instructional approach, and subject and content-like contextual factors demand, a wise choice for the students to improve attentiveness in class is not to use network and internet-connected electronic devices such as cell phones, smartwatches, laptops, tablets, etc., during lectures, which not only distract students but also negatively impact in students’ participation in classroom activities. Students are strongly recommended to turn off notifications during class and follow daily communication management strategies such as setting disconnection time, no screen time, communication preferred time, time for social media interaction, and decreasing internet addiction, etc. Mind mapping diagrams are free-flowing and helpful for brainstorming, finding linkage between ideas/concepts/topics, and generating ideas, keeping the main topic at the center. One may choose to prepare or recommend students to create concept maps as they are more structured, focused on connecting different concepts, and follow a hierarchy.
@@
Many other effective notetaking methods are also in practice such as the charting method, outline method, sentence method, flow method, boxing method, question-evidence-conclusion (QEC) method, split page method, writing on slides, bullet point method, list method, and rapid logging method [40-46]. The choice of notetaking method depends on many factors such as learning style, subject matter and topic, lecture structure and content delivery method, context, learner’s personal goal and preferences, available time to prepare notes, and need for mathematical, logical, and visual sophistication.
@@ III.1 Class and Lecture Design
The courses were designed based on the backward design of effective teaching, prioritizing the intended learning outcomes and course objectives over the chapters to be covered in the course, as well as the choice of instructional methods and assessment types. The backward design instructional method is based on aligning instruction, course content, pedagogy, and assessment with the learning objective, in contrast to the traditional teaching method of starting the learning process based on teacher-designed content, delivering lectures, assigning problems, and assessing students based on what was taught [47]. In the backward design method, one first determines the desired learning objectives and goals that students are expected to learn by the end of the day, chapter, and the course. Second, assessments are determined and planned to check whether students have learned what they were expected to learn. Third, learning activities and tasks are prepared that will be done in the classroom and beyond to prepare students for assessments. Algebra-based physics is primarily designed for students majoring in life sciences rather than students majoring in physical science. Students pursuing careers in life sciences are often less motivated and less prepared for physics courses, which require active, goal-oriented learning rather than passive, teacher-centered methods. To address this, we used a backward design approach in teaching these courses.
@@
The classes met twice a week for 75 minutes each day. Pedagogy includes lecturing, simulations, demonstrations, problem-solving, hands-on activities, think-pair-share, and class work as multiple choice questions. The classwork covered topics discussed at the beginning and end of the lecture, which helped ensure that students arrived on time and stayed for the whole class and boosted class attendance [13]. Students were strongly recommended to take a digital detox in the classroom, which minimizes digital distractions and helps minimize multi-tasking, which is considered one of the major learning pitfalls [48]. Students using electronic devices to take notes are prone to multitasking, but multitasking during learning impacts comprehension and knowledge retention as attention gets divided, focus decreases, the brain needs to pay the task-switching costs on refocusing, the brain receives cognitive load usually more than it can handle with no stress, and working memory is disrupted [49]. The assignments include online quizzes, primarily focused on concept development, and instructor-prepared homework assignments focused on problem-solving. The instructor used a liquid syllabus [50], which is a dynamic, web-based, interactive, and student-engaging syllabus that includes visual elements such as figures and videos, links to resources, and QR codes. The liquid syllabus has fixability in terms of dates, learning activities, and so on, to ensure the course objective is met by the end of the semester. The liquid syllabus, featuring more multimedia components, was adopted, inspired by the fact that students rarely read the syllabus once at the beginning of the semester and need to be made aware of the instructors’ expectations, requirements, and core objectives for the content and the course. A study shows that a liquid syllabus provides students with what they need to succeed in the course, increases equity, fosters a sense of belonging and respect through student-friendly language and visuals, and offers a humanizing introduction for the course [51]. Students must complete one prerequisite course before enrolling in Physics I. At Fayetteville State University (FSU), students must have a minimum passing grade in one of 4 math courses: college trigonometry, algebra and trigonometry, pre-calculus mathematics I, or pre-calculus mathematics II.
Table 1: FSU’s University catalog to assign letter grade for earned grade points on a scale of 0–100 and a GPA out of 4.0.
Points 90 – 100 80 – 89.9 70 – 79.9 60 – 69.9 ¡ 60
Grade A B C D F
GPA 4.0 3.0 2.0 1.0 0.0
Figure 2: A Pie chart showing point distribution in introductory physics course. Exams weighs 56% towards the final grade. Each mid term exam has 12% weight.
@@
The grading practices outlined in Table 1 were communicated at the beginning of the semester and a couple of times thereafter. The grade points are made student-trackable to ensure students understand the grading practices, can identify their strengths and weaknesses, and stay focused on their goals. Assignments were structured to (i) reduce the negative impact of late work, keeping students responsible for submitting assignments on time, (ii) give chances to implement feedback for improvement, and (iii) provide opportunities for students to reflect on and improve their performance, which motivated student in learning by doing and revising. We focused on reducing the basis of grading by emphasizing academic achievement and excellence, excluding non-academic components such as attendance, seminar participation, group discussion, extra work for points, etc. Students’ final grade was determined by aggregating their scores across assignments, with each assignment weighted according to its assigned weight. Online tests were provided with three possible attempts for each question, in which a link to the relevant section in the eTextbook and available resources was provided. A hint with insight into how to answer the questions was made visible after the first attempt, and the detailed solutions were provided a minute after the deadline. Numerical values were randomized for each student, and the order of answer choices was shuffled in each attempt. The desired number of significant figures was made visible for each question. Each student got feedback on each attempt. The best score out of the maximum three attempts was recorded. Students are encouraged to increase attentiveness, as we presented in Fig. 1, and be active physically and mentally in the learning environment rather than just being passive listeners.
Recognizing that elaborated rehearsal, implementation of feedback, repetition of tasks, and gaining overall skills are crucial, we assigned a lower weight to problem-solving homework assignments and online quizzes than terminal exams and the final comprehensive exam. For the same reason, the final exam had a slightly larger weight than each terminal exam. A schematic of point-distribution is presented in a pie-chart in Fig. 2. The pie chart is presented only for reference. The students’ grade presented in this study to analyze students’ performance is taken only from the third mid-term exams, which is the last mid-term before the final exam. The midterm exams were the instructor-prepared questionnaires. The problems were mostly problem-solving, analytical analysis, and critical thinking rather than knowledge transfer. The third mid exam (analyzed here) covered two chapters, namely (1) work and energy and (2) Impulse and momentum. Thus, the grade presented in this study is not the final grade that students received; rather, it is a 12% contribution towards their final grade.
III.2 Study Design